OpenStudy (anonymous):

m^(-5) = 1/32 Help please! Will fan and medal! @sangya21

OpenStudy (freckles):

\[m^{-5}=\frac{1}{m^5} \\ \frac{1}{m^5}=\frac{1}{32} \\ m^5=32 \] can you solve for m now?

OpenStudy (anonymous):

maybe. the only next step that pops in my head is 5(square root of) 32.... i suck at math.

OpenStudy (freckles):

ok well you could just raise both sides to the 1/5th power

OpenStudy (jhannybean):

hint: to get rid of a power you take the square root of that power.

OpenStudy (freckles):

oh and i think that is what you meant

OpenStudy (freckles):

\[(32)^\frac{1}{5}\] i wouldn't call it 5(square root of) 32 though :p

OpenStudy (freckles):

just the 5th root of 32 or 32 to the 1/5th

OpenStudy (freckles):

32 can actually be written as 2^some power

OpenStudy (anonymous):

@ frekles what you put is not what i meant omf doubleG Poland gone Tomboy thank you!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

* @freckles

OpenStudy (freckles):

lol i'm not familiar with that slang \[m^5=32 \\ (m^5)^\frac{1}{5}=(32)^\frac{1}{5} \\ m^{5 \cdot \frac{1}{5}}=(32)^\frac{1}{5} \\ m^1=(32)^\frac{1}{5} \\ m=(32)^\frac{1}{5}\]

OpenStudy (freckles):

now (32)^(1/5) can be simplified further

OpenStudy (freckles):

you know 32 is even so 2 is a factor of 32 32=2(16) 2 is also a factor of 16 since 16 is even 32=2(2)(8) blah blah blah 32=2(2)(2)(2)(2)

OpenStudy (freckles):

how many factors of 2 are in 32?

OpenStudy (anonymous):

i already have the answer. thx anyhow!!!

OpenStudy (freckles):

np